OpenStudy (anonymous):

a ball is thrown straight up with an initial speed of 30 m/s. Calculate the height and velocity of the ball after 4s

OpenStudy (anonymous):

For velocity we're gonna use this:\[v _{f}=v _{i}+a \times t\]

OpenStudy (anonymous):

v = 30+(-9.8)*4 v = -9.2 m/s Negative speed means that the ball is already on its way down after 4 seconds.

OpenStudy (anonymous):

For hight: \[y _{f}=y _{i}+v _{i} \times t + \frac{ a \times t ^{2} }{ 2 }\]

OpenStudy (anonymous):

y = 0 + 30*4-9.8*4^2/2 y = 41.6 m